# Gravitational lensing inside and outside of a marginally unstable photon sphere in a general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limits

### v1

@@ Abstract
It is believed that rays bent inside and outside photon spheres could affect partially the black hole shadow images by the Event Horizon Telescope and the rays near photon spheres would be detected by near-future space observations. The features of the rays near the photon spheres in not only black hole spacetimes but also exotic spacetimes would be important since one will need them to exclude black-hole mimickers. The deflection angles of the rays deflected by photon spheres diverge logarithmically and we can treat them by a strong-deflection-limit analysis. The error of the strong-deflection-limit analysis becomes large if antiphoton spheres exist in the spacetimes and the analysis breaks down when the photon spheres and the antiphoton spheres degenerate to form a marginally unstable photon sphere. This is because the deflection angles of the rays bent by the marginally unstable photon sphere diverge in powers. In this paper, we extend Eiroa, Romero, and Torres’s method to gravitational lensing of rays inside and outside of the marginally unstable photon sphere in a general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limits and we apply it to a Reissner-Nordström spacetime and a Hayward spacetime with the marginally unstable photon sphere. We have also confirmed that the deflection angles in the strong deflection limits by the method converge correctly to the deflection angle without approximations, while there are the mismatches of the coefficient of the power-divergent term of the deflection angles of the rays deflected just outside of the marginally unstable photon sphere in a semi-analytic calculation by the author previously.
@@ I Introduction
Gravitational lenses are often used to survey dark objects such as the black holes, extrasolar planets, and dark matters Schneider_Ehlers_Falco_1992 ; Schneider_Kochanek_Wambsganss_2006 . The gravitational lenses have been mainly studied in a weak gravitational field for a hundred years since we do not have enough technology to detect astrophysical phenomena in strong gravitational fields until the last decade. Rays deflected by the photon spheres of the black holes and their gravitational lensing effects in strong gravitational fields have been intermittently investigated Hagihara_1931 ; Darwin_1959 ; Atkinson_1965 ; Luminet_1979 ; Ohanian_1987 ; Nemiroff_1993 ; Frittelli_Kling_Newman_2000 ; Virbhadra_Ellis_2000 ; Bozza:2001xd ; Bozza:2002zj ; Perlick:2003vg ; Nandi:2006ds ; Virbhadra:2008ws ; Hioki:2009na ; Bozza_2010 ; Tsupko:2017rdo ; Nakao:2018knn ; Gralla:2019xty ; Okabayashi:2020apz ; Hsieh:2021scb ; Aratore:2021usi ; Tsupko:2022kwi ; Aratore:2024bro ; Tsupko:2025hhf . The effects of the photon spheres on the apparent shape of stars collapsing into the black holes also have been investigated Ames_1968 ; Synge:1966okc ; Yoshino:2019qsh ; Koga:2025dii . It is believed that the rays deflected just inside and outside photon spheres could affect partially the observed ring images of the supermassive black hole candidates at the center of our galaxy and the elliptic galaxy M87 by the Event Horizon Telescope Akiyama:2019cqa ; EventHorizonTelescope:2022wkp and the rays bent near the photon spheres would be detected by near-future space observations Lupsasca:2024xhq ; Johnson:2024ttr . It is important to investigate the features of the rays near the photon spheres in not only black hole spacetimes but also exotic compact objects such as wormholes Ellis:1973yv ; Chetouani:1984qdm ; Perlick:2003vg ; Muller:2004dq ; Nandi:2006ds ; Muller:2008zza ; Tsukamoto:2012xs ; Ohgami:2015nra ; Tsukamoto:2016qro ; Tsukamoto:2016zdu ; Shaikh:2018kfv ; Shaikh:2018oul ; Bronnikov:2018nub ; Shaikh:2019jfr ; Bronnikov:2021liv ; Maeda:2021wnl ; Olmo:2023lil ; Bugaev:2023mlc ; Zhang:2024sgs ; Xavier:2024iwr ; Tsukamoto:2024pid ; Bugaev:2024dte ; Solodukhin:2025opw and naked singularity Virbhadra:1998dy ; Virbhadra:2002ju ; Hioki:2009na ; Shaikh:2018lcc ; Shaikh:2019itn since one needs to exclude exotic compact objects or black-hole mimickers to reinforce claims that the observed objects must be supermassive black holes.
Gravitational lensing of rays reflected by just outside of a photon sphere in a Schwarzschild spacetime in a strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$, where $b$ and $b_{\mathrm{m}}$ an impact parameter and a critical impact parameter of the ray, by Bozza et al. Bozza:2001xd . The deflection angle of the rays which pass near outside of the photon sphere diverges logarithmically can be expressed by
@@
where $\bar{a}_{+}$ and $\bar{b}_{+}$ are a coefficient and a term which are determined by the metric of the spacetime with the photon sphere. Eiroa, Romero, and Torres suggested a method to obtain the coefficient $\bar{a}_{+}$ and the term $\bar{b}_{+}$ numerically and they applied it to a Reissner-Nordström black hole spacetime Eiroa:2002mk while Bozza made a semi-analytic formula in a general, static, spherically symmetric, and asymptotically-flat spacetime with the photon sphere to obtain the coefficient $\bar{a}_{+}$ and the term $\bar{b}_{+}$ and got $\bar{a}_{+}$ in an analytic form and $\bar{b}_{+}$ partially numerically in the Reissner-Nordström black hole spacetime Bozza:2002zj . The strong-deflection-limits analysis has been intensively studied and it has been extended over the last two decades  Bozza:2002af ; Eiroa:2002mk ; Petters:2002fa ; Eiroa:2003jf ; Bozza:2004kq ; Bozza:2005tg ; Bozza:2006sn ; Bozza:2006nm ; Iyer:2006cn ; Bozza:2007gt ; Tsukamoto:2016qro ; Ishihara:2016sfv ; Tsukamoto:2016oca ; Tsukamoto:2016zdu ; Tsukamoto:2016jzh ; Shaikh:2018oul ; Shaikh:2019itn ; Takizawa:2021gdp ; Zhang:2024sgs ; Igata:2025taz . Tsukamoto Tsukamoto:2016oca ; Tsukamoto:2016jzh investigated an alternative of Bozza’s method Bozza:2002zj , obtained $\bar{a}_{+}$ and $\bar{b}_{+}$ in analytic forms in the Reissner-Nordström spacetime and confirmed that they are consistent with $\bar{a}_{+}$ and $\bar{b}_{+}$ obtained by Eiroa, Romero, and Torres Eiroa:2002mk and by Bozza Bozza:2002zj .
@@
In this paper, we extend Eiroa, Romero, and Torres’s method Eiroa:2002mk to gravitational lensing of rays just inside and outside of the marginally unstable photon sphere in the general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ and we apply it to the Reissner-Nordström spacetime and a Hayward spacetime Hayward:2005gi with the marginally unstable photon sphere. We also confirm that mismatches in the coefficients $\bar{c}_{+}$ which were calculated by the author in the semi-analytic method previously Tsukamoto:2020iez , are modified in the new approach.
This paper is organized as follows. We review the deflection angle of rays in Sec. II, we investigate extended Eiroa, Romero, and Torres’s method to obtain the deflection angle of the rays inside and outside of the marginally unstable photon sphere in the strong deflection limits in the general, static, spherically symmetric, and asymptotically-flat spacetime and we apply it to the Reissner-Nordström spacetime and the Hayward spacetime with the marginally unstable photon spheres in Sec. III, and we investigate observable in gravitational lensing by using the deflection angles and a lens equation in Sec. IV. We conclude our results in Sec. V. We use a unit that a light speed and Newton’s constant are unity in this paper.
@@ II Deflection angle of rays
where denotes a differentiation with respect to $r$ and functions with the subtract m denotes the functions at $r=r_{\mathrm{m}}$. The circular light orbits are unstable (stable) if a condition
@@
From the wave number of a ray $k^{\mu}\equiv\dot{x}^{\mu}$, where $x^{\mu}$ is the coordinates and the dot denotes a differentiation with respect to an affine parameter along the trajectory of the ray, the trajectory of the ray is obtained as
@@
We consider that the ray, which comes from a spatial infinity, is reflected by a lensing object at the closest distance $r=r_{0}$ and is away to the spatial infinity. In this paper, we do not consider a case that the rays coming from the spatial infinity, passes through a wormhole throat, and leave to another spatial infinity if we consider wormhole spacetimes. At the closest distance $r=r_{0}$, due to
@@
where functions with the subtract 0 denote the functions at the closest distance $r=r_{0}$, we obtain, from Eq. (12),
@@
From Eq. (14), the impact parameter $b$ of the ray defined as $b\equiv L/E$, where $E\equiv-g_{\mu\nu}t^{\mu}k^{\nu}$ and $L\equiv g_{\mu\nu}\varphi^{\mu}k^{\nu}$ are the conserved energy and the angular momentum of the ray, respectively, is expressed as
@@
is satisfied so that the spacetime has the marginally unstable photon sphere. Under the assumptions, we obtain
@@
$V^{\prime\prime\prime}(r_{\mathrm{m}},b_{\mathrm{m}})=-\frac{b_{\mathrm{m}}D^{\prime\prime}_{\mathrm{m}}}{B_{\mathrm{m}}C_{\mathrm{m}}}=-\frac{D^{\prime\prime}_{\mathrm{m}}}{A_{\mathrm{m}}B_{\mathrm{m}}}<0,$ (24)
and the deflection angle of the ray bent outside and inside of the marginally unstable photon sphere in the strong deflection limits $r_{0}\rightarrow r_{\mathrm{m}}\pm 0$ or $b\rightarrow b_{\mathrm{m}}\pm 0$ will be expressed by,
@@
where $c_{\pm}$ is the coefficient and $\bar{d}_{\pm}$ is the term. From the relation
@@ III.1 Reissner-Nordström spacetime
In the Reissner-Nordström black hole spacetime, gravitational lensing of rays bent by outside of a photon sphere in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ was investigated by Eiroa, Romero, and Torres Eiroa:2002mk in a numerical way and by Bozza Bozza:2002zj in a semi-analytical way, and extended source effects Eiroa:2002mk , the time delay of rays Sereno:2003nd , retrolensing Eiroa:2003jf ; Tsukamoto:2016oca ; Tsukamoto:2016jzh , and higher order terms of the deflection angle Tsukamoto:2022tmm ; Tsukamoto:2022uoz ; Sasaki:2023rdf were studied. The images of light sources at near the photon sphere in the Reissner-Nordström black hole spacetime were investigated in Ref. Aratore:2024bro ; Tsupko:2025hhf by using a strong-deflection-limit analysis with arbitrary source distances Bozza:2007gt . In the Reissner-Nordström naked singularity spacetime, the size of the photon sphere or the shadows Zakharov:2014lqa and gravitational lensing of rays deflected inside and outside of a photon sphere in the strong deflection limit $b\rightarrow b_{\mathrm{m}}\pm 0$ were investigated in Refs. Shaikh:2019itn ; Tsukamoto:2021fsz ; Tsukamoto:2021lpm .
@@
numerically for the rays just outside of the marginally unstable photon sphere in the Reissner-Nordström spacetime.
From Eqs. (28) and (29), for the rays just outside of the marginally unstable photon sphere in the Reissner-Nordström spacetime, we get
@@
in numerical.
The deflection angle $\alpha$ is shown in Fig. 2 and the percent error of the deflection angle of Eq. (3.12) defined by
@@
is shown in Fig. 3.
@@ III.2 Hayward spacetime
$c_{+}=2^{11/6}3^{2/3}5^{5/4}\int^{\infty}_{0}\frac{dy}{\sqrt{432y+900y^{2}+625y^{3}}}\sim 7.771.$ (51)
but they do not consider the constants $\bar{d}_{+}$ and $\bar{d}_{-}$ and the coefficient $\bar{c}_{-}$. Tsukamoto Tsukamoto:2020iez studied the deflection angle of the rays reflected outside of a marginally unstable photon sphere in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ and got the coefficient as, by using the semi-analytic method,
@@
$\bar{d}_{+}=-4\sqrt{\frac{6}{5}}+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi\sim-5.62607.$ (53)
The percent errors of the deflection angle of Eq. (3.12) with the coefficients $\bar{c}_{+}$ and the constants $\bar{d}_{+}$ are plotted in Fig. 6.
@@
$\bar{d}_{+}\sim-5.60584$ (60)
in numerical for the Hayward spacetime with the marginally unstable photon sphere. From Eqs. (28) and (29), we get the numerical values of
$\bar{c}_{-}\sim 10.4155$ (61)
@@
$\bar{d}_{-}\sim-5.69675.$ (62)
The deflection angle $\alpha$ in the Hayward spacetime with the marginally unstable photon sphere is shown in Fig. 5 and the percent error of the deflection angle of Eq. (3.12) with the coefficients $\bar{c}_{+}$ and the constants $\bar{d}_{+}$ is shown in Fig. 6.
@@
Figure 6: The percent errors of the deflection angles $\alpha$ of Eq. (3.12) against the deflection angle $\alpha$ of Eq. (2.17) in the Hayward spacetime are plotted in top and bottom panels for the rays producing images outside and inside of the marginally unstable photon sphere, respectively. In the top pannel, the dashed (green), dotted (black), and solid (red) curves denote the percent errors of deflection angles $\alpha$ of Eq. (3.12) with $\bar{c}_{+}$ of Eq. (59) and $\bar{d}_{+}$ of Eq. (60) in our numerical method, with $\bar{c}_{+}$ of Eq. (50) and $\bar{d}_{+}=0$ by Chiba and Kimura Chiba:2017nml , and with $\bar{c}_{+}$ of Eq. (52) and $\bar{d}_{+}$ of Eq. (53) by Tsukamoto Tsukamoto:2020iez , respectively. In the bottom panel, the dashed (green) curve denote the percent error of the deflection angle $\alpha$ of Eq. (3.12) with $\bar{c}_{-}$ of Eq. (61), and $\bar{d}_{-}$ of Eq. (62) in our numerical method.
We have confirmed the coefficient $\bar{c}_{+}$ (50) by Chiba and Kimura Chiba:2017nml and the constant $\bar{d}_{+}$ (53) by Tsukamoto Tsukamoto:2020iez are correct while the coefficient $\bar{c}_{+}$ (52) by using the semi-analytic method in Ref.Tsukamoto:2020iez should be modified.

### v2

@@ Abstract
It is believed that rays bent inside and outside photon spheres could affect partially the black hole shadow images by the Event Horizon Telescope and the rays near photon spheres would be detected by near-future space observations. The investigation of the rays near the photon spheres in not only black hole spacetimes but also exotic spacetimes would be important since one will need them to exclude black-hole mimickers. The deflection angles of the rays deflected by the photon spheres diverge logarithmically and we can treat them by a strong-deflection-limit analysis. The error of the strong-deflection-limit analysis becomes large if antiphoton spheres exist in the spacetimes and the analysis breaks down when the photon spheres and the antiphoton spheres degenerate to form a marginally unstable photon sphere. This is because the deflection angles of the rays bent by the marginally unstable photon sphere diverge in powers. In this paper, we extend Eiroa, Romero, and Torres’s method to gravitational lensing of rays inside and outside of the marginally unstable photon sphere in a general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limits and we apply it to a Reissner-Nordström spacetime and a Hayward spacetime with the marginally unstable photon sphere. We have also confirmed that the deflection angles in the strong deflection limits by the method converge correctly to the deflection angle without approximations, while there are the mismatches of the coefficient of the power-divergent term of the deflection angles of the rays deflected just outside of the marginally unstable photon sphere in a semi-analytic calculation by the author previously.
@@ I Introduction
Gravitational lenses are often used to survey dark objects such as the black holes, extrasolar planets, and dark matters Schneider_Ehlers_Falco_1992 ; Schneider_Kochanek_Wambsganss_2006 . The gravitational lenses have been mainly studied in a weak gravitational field for a hundred years since we do not have enough technology to detect astrophysical phenomena in strong gravitational fields until the last decade. Rays deflected by the photon spheres of the black holes and their gravitational lensing effects in strong gravitational fields have been intermittently investigated Hagihara_1931 ; Darwin_1959 ; Atkinson_1965 ; Luminet_1979 ; Ohanian_1987 ; Nemiroff_1993 ; Frittelli_Kling_Newman_2000 ; Virbhadra_Ellis_2000 ; Bozza:2001xd ; Bozza:2002zj ; Perlick:2003vg ; Nandi:2006ds ; Virbhadra:2008ws ; Hioki:2009na ; Bozza_2010 ; Tsupko:2017rdo ; Nakao:2018knn ; Gralla:2019xty ; Okabayashi:2020apz ; Hsieh:2021scb ; Aratore:2021usi ; Tsupko:2022kwi ; Aratore:2024bro ; Feleppa:2024vdk ; Feleppa:2025ejh ; Tsupko:2025hhf ; Saketh:2025cwf . The effects of the photon spheres on the apparent shape of stars collapsing into the black holes also have been investigated Ames_1968 ; Synge:1966okc ; Yoshino:2019qsh ; Koga:2025dii .
It is believed that the rays deflected just inside and outside photon spheres could affect partially the observed ring images of the supermassive black hole candidates at the center of our galaxy and the elliptic galaxy M87 by the Event Horizon Telescope Akiyama:2019cqa ; EventHorizonTelescope:2022wkp and the rays bent near the photon spheres would be detected by near-future space observations Lupsasca:2024xhq ; Johnson:2024ttr . It is important to investigate the features of the rays near the photon spheres in not only black hole spacetimes but also exotic compact objects such as wormholes Ellis:1973yv ; Chetouani:1984qdm ; Perlick:2003vg ; Muller:2004dq ; Nandi:2006ds ; Muller:2008zza ; Tsukamoto:2012xs ; Ohgami:2015nra ; Tsukamoto:2016qro ; Tsukamoto:2016zdu ; Shaikh:2018kfv ; Shaikh:2018oul ; Bronnikov:2018nub ; Shaikh:2019jfr ; Bronnikov:2021liv ; Maeda:2021wnl ; Olmo:2023lil ; Bugaev:2023mlc ; Zhang:2024sgs ; Xavier:2024iwr ; Tsukamoto:2024pid ; Bugaev:2024dte ; Solodukhin:2025opw and naked singularity Virbhadra:1998dy ; Virbhadra:2002ju ; Hioki:2009na ; Shaikh:2018lcc ; Shaikh:2019itn since one needs to exclude exotic compact objects or black-hole mimickers to reinforce claims that the observed objects must be supermassive black holes.
Gravitational lensing of rays reflected by just outside of a photon sphere in a Schwarzschild spacetime in a strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$, where $b$ and $b_{\mathrm{m}}$ are the impact parameter and the critical impact parameter of the ray, by Bozza et al. Bozza:2001xd . The deflection angle of the rays which pass near outside of the photon sphere diverges logarithmically can be expressed by
@@
where $\bar{a}_{+}$ and $\bar{b}_{+}$ are a coefficient and a term which are determined by the metric of the spacetime with the photon sphere. Eiroa, Romero, and Torres suggested a method to obtain the coefficient $\bar{a}_{+}$ and the term $\bar{b}_{+}$ numerically and they applied it to a Reissner-Nordström black hole spacetime Eiroa:2002mk while Bozza made a semi-analytic formula in a general, static, spherically symmetric, and asymptotically-flat spacetime with the photon sphere to obtain the coefficient $\bar{a}_{+}$ and the term $\bar{b}_{+}$ and got $\bar{a}_{+}$ in an analytic form and $\bar{b}_{+}$ partially numerically in the Reissner-Nordström black hole spacetime Bozza:2002zj . The strong-deflection-limits analysis has been intensively studied and it has been extended over the last two decades  Bozza:2002af ; Eiroa:2002mk ; Petters:2002fa ; Eiroa:2003jf ; Bozza:2004kq ; Bozza:2005tg ; Bozza:2006sn ; Bozza:2006nm ; Iyer:2006cn ; Bozza:2007gt ; Tsukamoto:2016qro ; Ishihara:2016sfv ; Tsukamoto:2016oca ; Tsukamoto:2016zdu ; Tsukamoto:2016jzh ; Shaikh:2018oul ; Shaikh:2019itn ; Takizawa:2021gdp ; Zhang:2024sgs ; Igata:2025taz . Tsukamoto Tsukamoto:2016oca ; Tsukamoto:2016jzh investigated an alternative of Bozza’s method Bozza:2002zj , obtained $\bar{a}_{+}$ and $\bar{b}_{+}$ in analytic forms in the Reissner-Nordström black hole spacetime and confirmed that they are consistent with $\bar{a}_{+}$ and $\bar{b}_{+}$ obtained by Eiroa, Romero, and Torres Eiroa:2002mk and by Bozza Bozza:2002zj for the Reissner-Nordström black hole spacetime.
@@
In this paper, we extend Eiroa, Romero, and Torres’s method Eiroa:2002mk to gravitational lensing of rays just inside and outside of the marginally unstable photon sphere in the general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ and we apply it to the Reissner-Nordström spacetime and a Hayward spacetime Hayward:2005gi with the marginally unstable photon sphere. We also confirm that mismatches in the coefficients $\bar{c}_{+}$ which were calculated by Tsukamoto in the semi-analytic method previously Tsukamoto:2020iez , are modified in the new approach.
This paper is organized as follows. We review the deflection angle of rays in Sec. II, we investigate extended Eiroa, Romero, and Torres’s method to obtain the deflection angle of the rays inside and outside of the marginally unstable photon sphere in the strong deflection limits in the general, static, spherically symmetric, and asymptotically-flat spacetime and we apply it to the Reissner-Nordström spacetime and the Hayward spacetime with the marginally unstable photon spheres in Sec. III, and we investigate observable in gravitational lensing by using the deflection angles and a lens equation in Sec. IV. We conclude our results in Sec. V and discuss in Sec. VI. We use a unit that a light speed and Newton’s constant are unity in this paper.
@@ II Deflection angle of rays
where denotes a differentiation with respect to $r$ and functions with the subscript m denotes the functions at $r=r_{\mathrm{m}}$. The circular light orbits are unstable (stable) if a condition
@@
From
$k^{\mu}k_{\mu}=0,$ (12)
where $k^{\mu}$ is the wave number of a ray defined as $k^{\mu}\equiv\dot{x}^{\mu}$, where $x^{\mu}$ is the coordinates and the dot denotes a differentiation with respect to an affine parameter along the trajectory of the ray, the trajectory of the ray is obtained as
@@
We consider that the ray, which comes from a spatial infinity, is reflected by a lensing object at the closest distance $r=r_{0}$ and is away to the spatial infinity. In this paper, we do not consider a case that the rays coming from the spatial infinity, passes through a wormhole throat, and leave to another spatial infinity if we consider a wormhole spacetime. At the closest distance $r=r_{0}$, due to
@@
where functions with the subscript 0 denote the functions at the closest distance $r=r_{0}$, we obtain, from Eq. (13),
@@
From Eq. (15), the impact parameter $b$ of the ray defined as $b\equiv L/E$, where $E\equiv-g_{\mu\nu}t^{\mu}k^{\nu}$ and $L\equiv g_{\mu\nu}\varphi^{\mu}k^{\nu}$ are the conserved energy and angular momentum of the ray, respectively, is expressed as
@@
are satisfied so that the spacetime has the marginally unstable photon sphere. Under the assumptions, we obtain
@@
$V^{\prime\prime\prime}(r_{\mathrm{m}},b_{\mathrm{m}})=-\frac{b_{\mathrm{m}}^{2}D^{\prime\prime}_{\mathrm{m}}}{B_{\mathrm{m}}C_{\mathrm{m}}}=-\frac{D^{\prime\prime}_{\mathrm{m}}}{A_{\mathrm{m}}B_{\mathrm{m}}}<0,$ (25)
and the deflection angles of the rays bent outside and inside of the marginally unstable photon sphere in the strong deflection limits $r_{0}\rightarrow r_{\mathrm{m}}\pm 0$ or $b\rightarrow b_{\mathrm{m}}\pm 0$ will be expressed by,
@@
where $c_{\pm}$ are coefficients and $\bar{d}_{\pm}$ are finite terms. From the relation
@@ III.1 Reissner-Nordström spacetime
In the Reissner-Nordström black hole spacetime, gravitational lensing of rays bent by outside of a photon sphere in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ was investigated by Eiroa, Romero, and Torres Eiroa:2002mk in a numerical way and by Bozza Bozza:2002zj in a semi-analytical way, and extended source effects Eiroa:2002mk , the time delay of rays Sereno:2003nd , retrolensing Eiroa:2003jf ; Tsukamoto:2016oca ; Tsukamoto:2016jzh , and higher order terms of the deflection angle Tsukamoto:2022tmm ; Tsukamoto:2022uoz ; Sasaki:2023rdf were studied. The images of light sources at near the photon sphere in the Reissner-Nordström black hole spacetime were investigated in Ref. Aratore:2024bro ; Tsupko:2025hhf by using a strong-deflection-limit analysis with arbitrary source distances Bozza:2007gt . In the Reissner-Nordström naked singularity spacetime, the size of the photon sphere or the shadows Zakharov:2014lqa and gravitational lensing of rays deflected inside and outside of a photon sphere in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ were investigated in Refs. Shaikh:2019itn ; Tsukamoto:2021fsz ; Tsukamoto:2021lpm .
@@
for the rays just outside of the marginally unstable photon sphere, and
@@
for the ones inside of it in the Reissner-Nordström spacetime.
The deflection angles $\alpha$ are shown in Fig. 2 and the percent error of the deflection angle of Eq. (3.12) defined by
@@
is shown in Fig. 3. From Figs. 2 and 3, we confirm that our results are the almost same as ones by Sasaki Sasaki:2025web and that they converge to the deflection angle (2.17) in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$. At a glance at Fig. 2, it might seem as the deflection angle by Tsukamoto Tsukamoto:2020iez converge to the deflection angle (2.17) in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ but Fig. 3 shows that it does not converge as pointed out by Sasaki Sasaki:2025web .
@@ III.2 Hayward spacetime
$c_{+}=2^{11/6}3^{2/3}5^{5/4}\int^{\infty}_{0}\frac{\mathrm{d}y}{\sqrt{432y+900y^{2}+625y^{3}}}\sim 7.771.$ (52)
but they do not consider the constants $\bar{d}_{+}$ and $\bar{d}_{-}$ and the coefficient $\bar{c}_{-}$. Tsukamoto Tsukamoto:2020iez studied the deflection angle of the rays reflected outside of the marginally unstable photon sphere in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ and got the coefficient as, by using the semi-analytic method,
@@
$\bar{d}_{+}=-4\sqrt{\frac{6}{5}}+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi\sim-5.62607,$ (54)
where an integral
$I_{\mathrm{R}}(r_{\mathrm{m}})=2\int^{1}_{0}\left[\sqrt{\frac{6-3z+3z^{2}-z^{3}}{z^{3}(5-5z+z^{2})}}-\sqrt{\frac{6}{5z^{3}}}\right]\mathrm{d}z$ (55)
was calculated in numerical.
@@
$\bar{d}_{+}\sim-5.60958$ (62)
for the rays outside of the marginally unstable photon sphere
$\bar{c}_{-}\sim 10.4154$ (63)
@@
$\bar{d}_{-}\sim-5.67374.$ (64)
for the ones inside of the marginally unstable photon sphere in the Hayward spacetime.
The deflection angles $\alpha$ in the Hayward spacetime with the marginally unstable photon sphere are shown in Fig. 5 and the percent errors of the deflection angles of Eq. (3.12) with the coefficients $\bar{c}_{\pm}$ and the constants $\bar{d}_{\pm}$ are shown in Fig. 6. From Fig. 5, one may consider that the deflection angle by Tsukamoto Tsukamoto:2020iez converges to the deflection angle (2.17) in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ but Fig. 6 shows that it does not converge. From Fig. 6, we find that the deflection angle which we obtained in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ and one by Chiba and Kimura Chiba:2017nml in the strong deflection limit $b\rightarrow b_{\mathrm{m}}+0$ converges to the deflection angle (2.17) and we notice that the errors of our calculations are smaller than one by Chiba and Kimura Chiba:2017nml due to their lack of the constant $\bar{d}_{+}$. Therefore, we recognize that the coefficient $\bar{c}_{+}$ (51) by Chiba and Kimura Chiba:2017nml and the constant $\bar{d}_{+}$ (55) by Tsukamoto Tsukamoto:2020iez are correct while the coefficient $\bar{c}_{+}$ (53) by using the semi-analytic method in Ref.Tsukamoto:2020iez should be modified.
@@
Figure 6: The percent errors of the deflection angles $\alpha$ of Eq. (3.12) against the deflection angle $\alpha$ of Eq. (2.17) in the Hayward spacetime are plotted in top and bottom panels for the rays producing images outside and inside of the marginally unstable photon sphere, respectively. In the top pannel, the dashed (green), dotted (black), and solid (red) curves denote the percent errors of the deflection angles $\alpha$ of Eq. (3.12) with $\bar{c}_{+}$ of Eq. (61) and $\bar{d}_{+}$ of Eq. (62) in our numerical method, with $\bar{c}_{+}$ of Eq. (51) and $\bar{d}_{+}=0$ by Chiba and Kimura Chiba:2017nml , and with $\bar{c}_{+}$ of Eq. (53) and $\bar{d}_{+}$ of Eq. (55) by Tsukamoto Tsukamoto:2020iez , respectively. In the bottom panel, the dashed (green) curve denote the percent error of the deflection angle $\alpha$ of Eq. (3.12) with $\bar{c}_{-}$ of Eq. (63), and $\bar{d}_{-}$ of Eq. (64) in our numerical method.